### v1

# Automatically Detecting Checked-In Secrets in Android Apps: How Far Are We?

Abstract.
Mobile apps are predominantly integrated with cloud services to benefit from enhanced functionalities. Adopting authentication using secrets such as API keys is crucial to ensure secure mobile-cloud interactions. However, developers often overlook the proper storage of such secrets, opting to put them directly into their projects. These secrets are checked into the projects and can be easily extracted and exploited by malicious adversaries. While many researchers investigated the issue of checked-in secret in open-source projects, there is a notable research gap concerning checked-in secrets in Android apps deployed on platforms such as Google Play Store. Unlike open-source projects, the lack of direct access to the source code and the presence of obfuscation complicates the checked-in secret detection for Android apps. This motivates us to conduct an empirical analysis to measure and compare the performance of different checked-in secret detection tools on Android apps. We first conducted a literature review to find all the checked-in secret detection tools that can be applied to Android apps. Then, we evaluate three representative tools on 5,135 Android apps, comparing their performance and analyzing their limitations. Our experiment reveals 2,142 checked-in secrets affecting 2,115 Android apps. We also disclose that the current checked-in secret detection techniques suffer from key limitations.
All of the evaluated tools can miss a significant number of checked-in secrets in Android apps.
Nevertheless, we observed that the tools are complimentary, suggesting the possibility of developing a more effective checked-in secret detection tool by combining their insights. Additionally, we propose that analyzing string groups within methods containing checked-in secrets may provide a more effective strategy to overcome obfuscation challenges.
Android apps, Checked-in Secret, Empirical Study
ccs: Security and Privacy Detections of Leaked Checked-in Secrets
1. Introduction
Cloud computing is one of the most prevalent technologies in the modern software industry (Michalowski, [n. d.]).
Cloud services typically require clients to use secrets (e.g., API keys or API secrets) for authentication and secure access. However, in practice, developers often neglect the protection of the secrets and check them into their projects (Basak et al., 2023c).
For instance, the usage of hard-coded secrets is one of the 25 most dangerous software weaknesses (CWE, 2023). Such checked-in secrets can be exploited once extracted from the released software projects (798, 2024).
As an example, malicious users can use the extracted secrets to gain access to sensitive information. Recently, the database secrets of Uber company were mistakenly uploaded to GitHub, resulting in a leakage of information of 50,000 Uber drivers (ube, 2016).
Most of the existing studies focused on analyzing the checked-in secret problem in code-sharing platforms such as Github or Gitlab. For example, Meli et al. (Meli et al., 2019) designed a framework based on regular expressions that can identify potential checked-in secrets within a GitHub repository. Feng et al. (Feng et al., 2022) implemented a machine learning tool, PassFinder, to predict whether a string is a checked-in secret based on its contextual surroundings and intrinsic value. There also exist a number of community tools that can detect checked-in secrets in code-sharing platforms (tru, 2024; ggs, 2024; git, 2022; rep, 2024; whi, 2024; gsh, 2024). Other than developing a checked-in secret detection tool, some studies (Basak et al., 2023c; Rahman et al., 2022; Krause et al., 2022) conducted empirical analyses to investigate the checked-in secret issue on the code-sharing platform.
Meanwhile, there is a noticeable research gap of the checked-in secret problem on Android apps that are deployed on the Google Play Store. Android apps are also known to be susceptible to the problem of checked-in secrets (Basak et al., 2023c). Checked-in secrets can be extracted from the deployed Android apps through unpacking and reverse engineering (Renato, [n. d.]). Nevertheless, few studies have investigated this problem. Zuo et al. (Zuo et al., 2019) proposed the only tool, LeakScope, for checked-in secret detection in Android apps. Their study revealed that over 15,000 Android apps have the checked-in secret problem. However, their study is limited to three cloud providers, Google, Amazon, and Microsoft and requires substantial time and computational resources to expand its detection capabilities and analyze Android apps.
In this paper, we aim to evaluate the state-of-the-art checked-in secret detection tools when applied in Android apps and provide insights for the future development of Android checked-in secret detectors. We evaluate the performance of Android-specific checked-in secret detector, LeakScope, as well as adapting source-code-level checked-in secret detectors designed for code-sharing platforms. The source-code-level checked-in secret detectors can be adapted to Android apps since the source code of the apps can be recovered through decompiling or reverse engineering. Meanwhile, it is still challenging to directly leverage the source-code level checked-in secret detectors in Android apps due to the wide adoption of obfuscation. Obfuscation obscures the lexical information of the code that is often needed by the source-code-level checked-in secret detectors.
We aim to answer the following research questions in our paper:
  • RQ1: (Survey of existing tools) What are the existing checked-in secret detection tools? Which tools can be adapted to Android apps?
  • RQ2: (Evaluation and comparison) How does each secret detection tool perform when adapted to Android apps, and how do these tools compare against one another?
  • RQ3: (Implications) What are the limitations of the evaluated tools? What are the major implications of our research?
To address the proposed research questions, we conducted a systematic literature review to identify all academic and community checked-in secret detection tools. We discovered seven academic checked-in secret detection tools and six community tools. We further categorized the identified tools into three categories based on their methodology: Intrinsic value analysis, static analysis and machine learning-based analysis. We selected the framework proposed by Meli et al. (intrinsic value)  (Meli et al., 2019), LeakScope (static)  (Zuo et al., 2019), and PassFinder (machine learning)  (Feng et al., 2022) as the representative tool for each category.
We then made necessary modifications and extensions for each tool to be adapted to Android apps and launched them on 5,135 Android apps deployed on the Google Play Store. Overall, 2,142 valid checked-in secrets were detected in 2,115 Android apps. We then compared the performance of each tool based on the number of detected checked-in secrets and produced false positives/negatives. Our findings indicate that, although all the selected tools limited the presence of false positives, they still have significant limitations when applied to Android apps.
We observed that all the evaluated tools have a significant number of false negatives. The detection capability of both Meli et al.’s work and LeakScope is largely constrained by pre-defined inputs to specify the detectable types of checked-in secrets (i.e., regular expressions for Meli et al.’s work and cloud API signatures for LeakScope). In contrast, PassFinder’s detection capability is not limited by pre-defined inputs, but the presence of obfuscation in Android apps greatly reduces its performance.
Despite the limitations of each tool, we observe that these approaches can complement each other. For instance, the data flow analysis of checked-in secrets performed by LeakScope could assist Meli et al.’s work in extracting checked-in secrets that would otherwise go undetected. The methodology utilized by Meli et al. can enhance checked-in secret prediction accuracy as well. Additionally, analyzing string groups within methods containing checked-in secrets could be a more effective indicator of the presence of checked-in secrets, as obfuscation does not alter these strings.
To summarize, we make the following major contributions in this paper:
  • To the best of our knowledge, we are the first to conduct an empirical study of the checked-in secret issues in Android apps deployed on the Google Play Store.
  • We adapted existing checked-in secret detection tools that are originally designed for open-source projects to Android apps deployed on the Google Play Store.
  • We evaluated and compared the performance of the existing checked-in secret detection tools. We analyzed the limitations of existing checked-in secret detection tools and proposed future research directions for developing more effective checked-in secret detection tools tailored for Android apps.
  • We provided a benchmark dataset consisting of the code snippets of the validated checked-in secrets in Android apps.
Data availability. Our code and dataset are available on our project website (dat, [n. d.]). For the dataset, we provided sanitized code snippets and data flow of the detected checked-in secrets. Due to the sensitiveness of the dataset, we did not provide the original Android apps we used in the experiment. The Android apps can be shared to selected researchers and developers upon request.
2. Background
2.1. Checked-in Secrets
Cloud service providers deploy various methods for authenticating processes to minimize security risks (Lu, 2011). Using the secrets to authenticate cloud service is the most common approach. When incorporating cloud services into an application, cloud service providers require developers to generate secrets.
These secrets allow the cloud service providers to identify which project is calling the cloud API, verify its access to certain resources, and manage billing functions.
Leaking the sensitive cloud service secrets to the public can result in severe security consequences for both cloud service providers and applications. Access to the checked-in secrets allows anyone to utilize the allocated resources and functionalities. It can lead to (1) data leakage if the checked-in secret allows for access to application data, (2) data corruption if the checked-in secret allows for data manipulation, (3) monetary losses if it creates costs for invoking the concerned APIs, and (4) denial of service (DoS) if there is a limit for the API invocation.
{listing}
Code snippet with one checked-in secret. It is extracted from an Android app in our experiments.
{minted}
[linenos, frame=single, obeytabs=true, numbersep=-10pt]java task.addOnSuccessListener(listener) context = r0.getApplicationContext()¿() r7.ggh(context, ”AIza——”) r0.setupLifecycleListener()
As an example, listing 1 is a code snippet extracted from one Android app. The string at line 3 that starts with ”Alza” is a Google API key, which could be used to grant access to Google Map services. This secret can be used to perform a DoS attack since the API has a limit of requests per minute or cause financial losses since it costs \$0.005 per request.
2.2. APK Files
Unlike the open-source Android app projects on GitHub, the Android apps deployed on the Google Play Store are distributed as APK (Android Package Kit) files. An APK file is essentially an archive file that contains all the information needed for an Android app to be successfully executed on a device. When a user wants to download an Android app from platforms such as the Google Play Store, the platforms will automatically download the APK files for the users. Several automatic tools exist, such as ApkTool (apk, 2024), dex2jar (dex, 2024), Soot (soo, 2024) and Mobsf (mob, 2024), that can perform reverse engineering on an APK file. However, these reverse engineering tools all have limitations. They may not accurately recover the source code from APK files. Sometimes, they might even miss some classes while analyzing the APK files. Furthermore, APK files commonly adopt obfuscation techniques to enhance the application’s security (Dong et al., 2018). Obfuscation obscures the code of the Android apps by altering the names and hierarchies of code artifacts. For example, in line 3 of Listing  4.1.4, the method ggh was originally a method called ”initialize()” provided by Google to send an API request. Those factors are the main challenges that need to be addressed for the tools focused on Android apps. The effectiveness of detection tools relying on code context is significantly reduced by obfuscation, which makes variables and class names meaningless.
3. RQ1: Survey of Existing Tools
The checked-in secrets issue is a long-studied subject in the software development field. There exist many tools that aim to detect checked-in secret. In RQ1, we aim to find all these available academic and community checked-in secret detection tools and analyze whether they can be adopted on Android apps through a detailed literature review process.
3.1. RQ1-Methodology: Literature Review
We conducted a literature review to gather a comprehensive list of checked-in secret detection tools/methodologies. Our search query is composed of two parts, as shown in Table 1: Part 1 includes terms related to secret types (e.g., API key, secret, password), while Part 2 focuses on detection methodologies.
Table 1. Search String Group
Part 1 Part 2
(API) key, secret, credential, password detection, leakage, exposure
Each search query is a combination of Part 1 and Part 2, resulting in a total of 12 queries (i.e., 4*3), such as ”API key detection.” We initially searched these keywords across Google Scholar, IEEE Xplore, and the ACM Digital Library.
Each query returns a large number of search results, including irrelevant ones. For example, some queries, such as API key detection, returned over 10,000 results on ACM digital library. Manually analyzing and selecting relevant papers from such a large pool is impractical. Therefore, for each search query, only the first 500 results were analyzed. In total, 5,500 papers were selected for our initial filtering process.
During the initial filtering process, any paper with a title not related to terms like API, security, secret detection, or privacy leakage was excluded. This step resulted in the selection of 78 papers. Subsequently, two authors conducted a more thorough review of these 78 papers to identify those most relevant to our subject. In this stage, both authors read the abstracts (and, if necessary, the full content) of each paper to determine its relevance. If there was any disagreement, a discussion was held to reach a consensus. 19 papers were selected after this step.
We also conducted a snowballing process, a strategy that is commonly adopted in other literature review research (Liu et al., 2022). This ensures we do not miss any important papers on the checked-in secret. For each paper that we initially selected, we reviewed references from the selected papers as well as papers that cited them. This process continued until no new relevant papers were identified.
Eventually, we found 29 academic papers that are related to checked-in secret detection.
For community tools, we used the six community tools selected by Setu et al.  (Basak et al., 2023a) as they conducted a thorough selection process to perform an analysis of checked-in secret detection tools on the open-source repositories.
3.2. RQ1-Result
Among the 29 papers, seven of them developed a checked-in secret detection tool or methodology of our interest, while others are either empirical studies on checked-in secret issue and management or only specific to platforms such as DockerHub. Overall, a total of 13 checked-in secrets detection tools when combined with community-developed tools.
We categorized the checked-in secret detection tools into three groups: Intrinsic value analysis, static analysis and Machine Learning based analysis according to their detection approach. Table 2 presents the list of all the identified checked-in secret detection tools and their corresponding categorization. Overall, there are seven Intrinsic value analysis tools, one static analysis tool and five Machine Learning based analysis tools.
Intrinsic value analysis. This type of checked-in secret detection tools focused on the intrinsic value of the strings under analysis. The analysis is generally done using two processes. First, regular expression matching. In the cases of certain cloud providers, the generated checked-in secrets often adhere to the same format. For instance, Google’s checked-in secrets always take the form of “ALza******”. Utilizing regular expression matching enables the extraction of such checked-in secrets from the code. Second, entropy calculation. As checked-in secrets are all randomly generated strings, the randomness of the characters in a string will be a good indicator of checked-in secret. These techniques leveraged Shannon entropy to give a good estimation of randomness. A randomly generated string will generally have higher entropy than a normal string.
Static analysis.
To the best of our knowledge, there has only been one tool that extracts checked-in secrets through static analysis, which is LeakScope(Zuo et al., 2019). By providing a location where a checked-in secret is used, static analysis can calculate the value of the checked-in secret through data and control flow analysis.
Machine Learning based analysis. This type of checked-in secret detection tool takes into consideration the fact that the contextual surroundings of checked-in secrets follow some specific patterns. Factors such as file type, class name, variable name and code snippet could be used as features to train a machine learning model to predict whether a string is a checked-in secret.
Table 2. Categorization and Tools
Category Tools Seletced
Intrinsic Value Analysis Meli et al (Meli et al., 2019)
Gitleak (git, 2022)
Repo Supervisor (rep, 2024)
Trufflehog (tru, 2024)
Whispers (whi, 2024)
GGshield (ggs, 2024)
Github Secret Scanner (gsh, 2024)
Static Analysis LeakScope (Zuo et al., 2019)
Machine Learning Analysis PassFinder (Feng et al., 2022)
Secrethunter (Wen et al., 2022)
Seagull  (Han et al., 2024)
Saha et al (Saha et al., 2020)
Lounici et al (Lounici et al., 2021)
3.3. Tool Selection for Further Evaluation.
The checked-in secret detection tools within the same category adopted similar techniques. For example, checked-in secret detection tools of type intrinsic value analysis either adopted regular expression matching or entropy calculation. Some tools adopted both of the two techniques. Therefore, we decided to select one representative tool for each category for further evaluation in RQ2 and RQ3.
For intrinsic value analysis, we chose the Three-Layer Filter technique proposed by Meli et al. (Meli et al., 2019) since (1) intrinsic value analysis highly relies on manually crafted regular expressions; the Three-Layer Filters provided a comprehensive list of precise regular expressions, and (2) the Three-Layer Filter is proven to be one of the best tools in this category (Basak et al., 2023a; Meli et al., 2019).
For the representative static analysis tool, we chose LeakScope because it is the only static analysis tool designed for Android APK files and is available on GitHub (Lea, 2020).
For machine learning-based tools, we selected PassFinder for further evaluation. PassFinder, SecretHunter and Seagull are the only publicly available tools. Among them, SecretHunter aims to address the issue of missing files in GitHub repositories during the scanning process, often due to API and bandwidth limitations. Since SecretHunter is tailored specifically for the GitHub platform, we excluded it from our study. As for Seagull, it highly relies on the results provided by CodeQL (cod, 2024), , which is incompatible with our focus on Android apps. Furthermore, the key idea of Seagull is very similar to PassFinder, which takes contextual information (i.e., the surrounding code snippets) as input for the models. Given their similarity in leveraging context for detection, we selected PassFinder as the representative tool for machine-learning-based analysis in this study.
By selecting these representative tools, we ensured that each category of checked-in secret detection methodology was evaluated in the context of Android apps. Below, we describe the three representative checked-in secret detection tools in detail.
Three-Layer Filter. Meli et al. (Meli et al., 2019) is the first to conduct a large-scale analysis on the checked-in secret problem of the open-source projects on GitHub, covering 13% of all the public repositories. Their methodology involved initially extracting strings from the repositories through regular expression matching. Subsequently, these candidate strings go through a Three-Layer Filter by evaluating the candidate string’s entropy value, pattern, and word. First, the authors designed 15 regular expressions for 12 different cloud providers and applied the 15 regular expressions to extract candidate strings from the source code. Then, the extracted candidate strings will be filtered by three layers. We refer to this technique as Three-Layer Filter for the rest of this paper. Specifically, the three filers are:
- Entropy Filter. The entropy filter would first calculate the Shannon Entropy (Shannon, 1951) for each string, and within the same regular expression group, a string will fail if its Shannon entropy is three standard deviations away from the mean.
- Word filter. Meli et al. also observed that if a string contains an English word, it is nearly impossible for the string to be a checked-in secret. Therefore, a string will fail if it contains words from a pre-defined set of English words.
- Pattern filter. Meli et al. also observed that there are some patterns that are unlikely to appear in checked-in secrets. Therefore, they proposed that a string should not be considered as checked-in secret if it contains patterns such as repetitive, ascending, and descending characters.
LeakScope (Zuo et al., 2019) is the first automated tool that aims to detect the data leakage problem caused by checked-in secret leakage and misconfiguration of secret permission in Android apps. Their tool is developed based on the key finding that if a project wants to use the services provided by cloud providers, the project would likely call the cloud APIs provided by the cloud providers (i.e., the APIs defined in the cloud SDKs). These cloud APIs generally take one or several checked-in secrets as part of their parameters. Take the CloudStorageAccount.parse(String) API of Microsoft Azure as an example; this cloud API takes a checked-in secret as a parameter and returns the account that is connected to this checked-in secret (clo, 2024); LeakScope collected 32 cloud APIs that take checked-in secret as parameter from three cloud providers: Microsoft Azure, Amazon AWS and Google Firebase. To extract the value of the checked-in secret, LeakScope performs an inter-procedural backward analysis starting from the cloud APIs; along with the backward analysis, LeakScope keeps track of the data flow associated with the parameter of this cloud API. Through this data flow tracking, LeakScope proceeds to compute the value of the parameter. Moreover, LeakScope employs a signature generation mechanism for cloud APIs based on their package structure and content, enabling identification of their usage within the project. This signature empowers LeakScope to operate effectively, even in scenarios where APK files are obfuscated.
PassFinder (Feng et al., 2022) is a checked-in secret detection tool utilizing the deep neural network. It conducts an analysis of strings by considering both the intrinsic value of the string itself and its contextual surroundings, particularly encompassing twelve lines of context. PassFinder trains separate models for the intrinsic value and context of the string. The prediction is given by combining the results of two models. PassFinder is designed to analyze the password and API credentials in the GitHub repository as it is trained with the code snippet collected from the GitHub repository. It is also capable of handling different programming languages.
4. RQ2: Comparsion of Exisiting TOOLs
In RQ2, we aim to evaluate and compare the performance of each category’s selected representative checked-in secret detection tools on Android apps.
4.1. RQ2-Methodology
We aim to compare the performance of the selected tools by answering the following two sub-research questions:
  • RQ2.1 (Number of detected checked-in secrets): How many checked-in secrets can be detected by each of the selected tools?
  • RQ2.2 (False positives): How many checked-in secrets reported by each tool are false positives?
  • RQ2.3 (False negatives): How many false negatives are produced by each tool?
4.1.1. Dataset
We first collected a dataset of Android apps as the evaluation subjects. We leveraged Raccoon APK Downloader (rac, 2024) to download Android apps from Google Play Store. We selected the top 200 Android apps in each category based on the number of downloads according to the statistics on AppBrain (app, 2024). After excluding the Android apps that failed to be processed by Soot, we collected 5,135 APK files in total.
The experiment was conducted on a High-Performance Computing cluster that consisted of 24 compute nodes with 32 cores and 512 GB of memory and approximately 10 TB of volatile-scratch disk space. We utilized a total of 32 CPUs for 16 jobs, with each job using 2 CPUs running concurrently.
4.1.2. Tool Adaptation
None of our selected tools are ready to use for our experiments. The Three-Layer Filter did not provide the necessary code, LeakScope did not provide the cloud API signatures, and PassFinder did not provide the pre-trained models. Thus, some implementation efforts were necessary for our experiments. Additionally, unlike PassFinder, whose detection capability is not strictly constrained, LeakScope and the Three-Layer Filter can only detected the checked-in secrets that are used as inputs in the pre-defined cloud APIs or adhere to the pre-defined regular expressions. Therefore, expanding the list of cloud APIs and regular expressions was necessary to equalize the detection capabilities of LeakScope and the Three-Layer Filter and conduct a fair comparison.
In this section, we detail the adaptations we made for all the tools.
Adaptation for Three-Layer Filter. As the source code for the filtering process of the Three-Layer Filter is not provided by the authors. We first implemented the necessary code for the filtering process following the explanation of author (Meli et al., 2019). For the string extraction, the source code is not directly accessible in APK files, we used Soot (soo, 2024), which is capable of identifying all the variables of type java.lang.String to extract all the strings within the APK files.
The authors provide regular expressions. However, we noticed that in their experiment, their filtering process only excluded 2% of the strings due to their precise regular expressions. Their regular expression rarely mismatch checked-in secrets. As a result, the filtering process has minimal impact on reducing false positives in their experiment.
However, it is challenging to obtain precise regular expressions for all the categories of checked-in secrets in practice. It is more common for cloud providers to adopt loose regular expressions. For example, checked-in secrets of Twitter and Facebook do not exhibit a strong pattern. Any combination of letters and numbers that fall within a specific range of lengths can match the regular expressions. Loose regular expressions would match more non-secret strings compared to precise regular expressions. Passing more non-secret strings to the filtering process is necessary to fully validate the performance of the Three-Layer Filter.
Therefore, we derived two more regular expressions, as shown in Table 3. We chose Twitter client ID and secret because the corresponding checked-in secrets do not exhibit a strong pattern; Twitter APIs are commonly used in Android apps.
Furthermore, the authors do not provide the source code for the Three-Layer Filter. Therefore, we also implemented the necessary code based on the logic proposed by the authors. The implementation and the full list of collected regular expressions can be found on our website (dat, [n. d.]).
Table 3. Newly proposed regular expressions
Cloud providers Regular expression
Twitter Client ID [0-9a-zA-Z]18,25
Twitter Client secret [0-9a-zA-Z]40,50
Adaptation for LeakScope. In order to use LeakScope to extract the checked-in secrets from an APK file, we have to provide the complete signature of the cloud API and specify which API parameter holds checked-in secrets to LeakScope. The authors provided the names of 32 cloud APIs from Amazon AWS, Microsoft Azure, and Google Firebase, but they did not provide the complete signatures of cloud APIs. Therefore, we also collected the detailed signatures of these 32 cloud APIs. The full list of signatures can be found on our website (dat, [n. d.]).
However, the provided cloud APIs are limited to three cloud providers. The number of supported cloud providers is very limited compared to the Three-Layer Filter. To perform a more in-depth and fair analysis, we aim to collect more cloud APIs that can be analysis starting points for LeakScope, i.e., the cloud APIs that take checked-in secret as a parameter. Given the abundance of cloud providers in the contemporary market, attempting to collect all the cloud APIs that take checked-in secret as parameter from every single provider would be infeasible, especially since some cloud providers may not offer comprehensive API documentation. Therefore, instead of solely relying on the API documentation provided by every cloud provider, our cloud API collection strategy also takes into account the potential cloud API extracted from our dataset.
We first collected 15 cloud providers, including those covered by the Three-Layer Filter but not by LeakScope, as well as the top cloud service providers in 2024 (Slingerland, [n. d.]). Then, we read the available API documentation of the collected cloud providers to identify the cloud APIs that use secrets as part of their parameters. To mitigate the risk of missing some relevant API documentation, we also adopted an additional step. We collected information of all the methods that are called at least once in the Android apps of our dataset. We then selected all the methods that contain the names of our collected cloud providers. By analyzing the context and implementation (API documentation if available) of these methods, we determined whether they are our targeted cloud APIs. For example, a constructor, OSSStsTokenCredentialProvider(String, String, String) was not originally found by reading the API documentation of Alibaba, but as its package name contains Alibaba, we were able to identify this cloud API.
Overall, we collected 13 new cloud APIs from five additional providers. For the remaining 10 cloud providers, we failed to recover any cloud APIs that take checked-in secrets as input. This results in a total of 48 cloud APIs from 10 cloud providers. Table 4 summarizes the distribution of these APIs across the different cloud providers. The authors of LeakScope provided 32 cloud APIs from Amazon, Google and Microsoft. By analyzing API documentation, we identified 10 additional cloud APIs from Alibaba, Facebook, Redhat, Twitter and Stripe. Furthermore, we collected 6 new cloud APIs from Alibaba, Facebook, Paypal and IBM Cloud.
We acknowledge that our collection process cannot capture every cloud API using checked-in secrets as parameters, but our selection is representative of the typical APIs used in Android apps. Overall, they greatly enhanced the detection capability of LeakScope.
Table 4. Collected cloud APIs and cloud providers’ information that are used as the parameters for LeakScope, the signature of each cloud API can be found on our website (dat, [n. d.]).
Cloud providers Number of Original provided Cloud APIs by LeakScope Number of Newly Collected Cloud APIs by API Documentation Number of Newly Collected Cloud APIs by API Extraction Total Number of Collected Cloud APIs
Amazon AWS 26 0 0 26
Alibaba Cloud 0 2 2 4
Microsoft Azure 4 0 0 4
Facebook/Meta 0 2 1 3
Twitter/X 0 2 0 2
Redhat 0 2 0 2
IBM Cloud 0 0 2 2
Google 2 0 0 2
Paypal Braintree 0 0 1 1
Stripe 0 2 0 2
Adaptation for PassFinder. We contacted the author to request the pre-trained models, but they indicated that the project is not currently being maintained, but the necessary dataset and instruction are provided (pas, 2023). As a result, we followed the exact instructions provided by the author to train the two models provided by PassFinder.
For the password model, we leveraged the original training set provided by the authors (pas, 2023). For the context model, the original training set contains 12,000 code snippets for the ten most popular programming languages on GitHub. For each language, there are 600 snippets with checked-in secrets and 600 snippets without checked-in secrets.
Once the password and context models were trained, we utilized the password model to identify potential checked-in secrets within our dataset. Subsequently, we employed Soot (soo, 2024) to extract the contextual surroundings of the identified checked-in secrets, which were then analyzed by the context model for detection. However, instead of using source code, Soot will output the contents of APK files in Jimple format, an intermediate representation of the Java language. Jimple representations are very similar to Java syntax, so we decided to remain with the Jimple format for the extracted contextual surroundings.
Table 5. The type of checked-in secrets that each tool can detect
Tools Detection capability (Cloud providers)
The Three-Layer Filter Amazon, Microsoft, Facebook, Twitter client ID, Twitter client secret, Twitter access token, Google API key, Stripe, Square, Picatic, PayPal Braintree, Twilio, MailGun, MailChimp
LeakScope Amazon, Microsoft, Facebook, Twitter client ID, Twitter client secret, Google API key, Stripe, Square Redhat, IBM cloud
PassFinder Not strictly constrained
4.1.3. Comparison of the Detection Capability of Each Tool
After making the necessary adaptation for each tool, the detection capability of each tool is listed in Table 5. Overall, we extended LeakScope’s detection capability to align with the Three-Layer Filter. However, it is infeasible to identify a comprehensive set of cloud API signatures for LeakScope that matched all the regular expressions used by the Three-Layer Filter and vice versa. This difficulty arises because some cloud providers do not issue secrets that follow specific patterns that can be precisely captured by regular expressions (e.g., Alibaba, Redhat and IBM). Furthermore, we failed to find the cloud APIs of the cloud providers that the Three-Layer Filter covered (e.g., MailGun, MailChimp, and Picatic). Although the detection capability of LeakScope and the Three-Layer Filter is not entirely equalized, our extension ensured a more meaningful comparison among the tools. In the following questions, we also show the detailed detection results of each cloud provider for a fair comparison.
4.1.4. Checked-in Secret Validation and Ethical Considerations.
The common challenge for evaluating the checked-in secret detection tool is the lack of the ground truth. To the best of our knowledge, there is not an available benchmark dataset for checked-in secrets in Android apps. A complete validation of a checked-in secret requires us to access the services associated with the checked-in secret, which would raise ethical concerns. As a result, we decide whether a checked-in secret is valid mainly by manually analyzing its format and contextual surroundings, which is a commonly adopted technique for checked-in secret related research  (Basak et al., 2023a, b). For example, in Listing 2, the presence of some functions from Google Cloud Library is an indicator of checked-in secret. Furthermore, two authors were responsible for manually verifying the potential checked-in secrets. In cases of disagreement, discussions were held where each author presented their reasoning and attempted to reach a consensus through mutual persuasion.
{listing}
Code snippet with one checked-in secret and its context. It is extracted from an Android app in our experiments.
{minted}
[linenos, frame=single, obeytabs=true, numbersep=-10pt]java GoogleApiClient GoogleClient = x.GoogleClient; d.b0.c.l.c(GoogleClient); mGoogleApiClient.disconnect(); GoogleSignInOptions.Builder builder = new Builder; builder = builder.requestIdToken(””); builder = builder.requestEmail();
In addition to manual analysis, we tested the validity of checked-in secrets categorized as Google API keys by sending requests to the Google Map service as they are not related to any form of sensitive information. We also ensured that the request sent by us would not cause any significant financial loss; each checked-in secret was only used once. Furthermore, for any detected checked-in secret, we notified the developers of the corresponding Android apps.
4.2. RQ2.1 Number of Detected Checked-in Secrets
4.2.1. RQ2.1-Methodology
In this research question, we aim to evaluate the effectiveness of the tools by evaluating the number of checked-in secrets that can be detected by each tool. As the detection capability of each tool is not identical, we evaluated the number of detected checked-in secrets identified by each tool across various cloud providers to provide a fairer comparison among the tools. The detailed result of the number of detected checked-in secrets is presented in Table 6. Note that the detected checked-in secrets in this table are not necessarily true positives. In this research question, we demonstrate the overall results reported by each tool. We conduct an analysis on the false positives of the tools in Section 4.3.
Table 6. Number of detected checked-in secrets corresponding to the cloud providers of the three selected tools, NA signifies that the checked-in secrets of the corresponding cloud providers are not in the detection capability of the tool.
Cloud Providers Three-Layer Filter LeakScope PassFinder Total # of Unique Checked-in Secrets per Cloud Provider
Amazon 0 5 0 5
Microsoft 0 0 0 0
Facebook 0 10 0 10
Twitter Access Token 0 NA 0 0
Twitter Client ID/Secret 88,172 10 0 88,182
Google 664 1,816 0 2,117
Stripe 0 0 0 0
Square 0 0 0 0
Picatic 0 NA 0 0
PayPal Braintree 0 0 0 0
Twilio 2 NA 0 2
MailGun 0 NA 0 0
MailChimp 0 NA 0 0
Redhat NA 0 0 0
IBM Cloud NA 0 0 0
Alibaba NA 0 0 0
Total # of Checked-in Secrets of Each Tool 88,838 1841 0 90,316
4.2.2. RQ2.1-Results
Table 6 lists the number of detected checked-in secrets of each tool. In total, 90,316 unique checked-in secrets were identified by the three tools. The Three-Layer Filter successfully identified 88,838 checked-in secrets. LeakScope successfully identified 1,841 checked-in secrets. For PassFinder, the password model successfully identified 11,343 checked-in secrets, but the context model failed to identify any checked-in secrets.
The Three-Layer Filter detected the highest number of unique checked-in secrets, while PassFinder failed to detect any. PassFinder’s limited effectiveness can be partially attributed to the context model’s underperformance, which was likely impacted by obfuscation. We further conducted a study to investigate the reasons behind such underperformance of PassFinder. Details of this study is discussed in Section 4.5.
We further categorized all the detected checked-in secrets of each tool to the corresponding cloud providers, as shown in Table 6. Among all the cloud service providers that are supported by each tool, only a few are detected in our dataset. Most of the detected checked-in secrets are from Google cloud service and Twitter cloud service, which is expected given Google’s status as one of the most popular cloud providers. Additionally, LeakScope was able to detect a greater number of Google checked-in secrets compared to the Three-Layer Filter.
For Twitter Client ID/Secret. The Three-Layer Filter identified 88,172, whereas LeakScope only identified 10. This is unsurprising as we adopted loose regular expressions for Twitter when applying the Three-Layer Filter. However, on analyzing a sample of the 88,172 checked-in secrets, despite that they are indeed random strings that appear to be checked-in secrets, none of them are truly Twitter checked-in secrets. This issue will be further discussed in RQ2.2.
Summary: Overall, there are more checked-in secrets that the Three-Layer Filter can identify compared to other tools, demonstrating its ability to detect random secret-like strings. However, this result alone does not demonstrate its effectiveness. Evaluating false positives and negatives is equally critical in assessing the tool’s effectiveness.
4.3. RQ2.2-False Positive
4.3.1. RQ2.2-Methodology
In this research question, we aim to compare the false positives reported by each tool. As described in Section 4.1.4, it is impractical and impossible to fully validate all the obtained checked-in secrets. Consequently, we conducted a manual analysis to identify the false positives. Furthermore, since only checked-in secrets belongs to Google API key and Twitter client ID/Secret are detected by at least two tools. We only selected these checked-in secrets for the comparison of false positives. Additionally, PassFinder is excluded from this comparison as it did not detect any checked-in secrets. A further experiment to comprehensively evaluate PassFinder is detailed in Section 4.5.
However, since the Three-Layer Filter detected 88,172 Twitter client IDs/Secrets, manually validating all of these checked-in secrets was unfeasible. To address this, we randomly sampled 383 checked-in secrets from the 88,172, ensuring a 95% confidence level with a 5% margin of error (ran, 2024). Then, we manually analyzed the randomly sampled Twitter client IDs/Secrets and all the Google API keys to identify any false positive as described in Section 4.1.4.
4.3.2. RQ2.2-Result
After manually analyzing the checked-in secrets of Google and Twitter, we discovered that for Google, Neither LeakScope or the Three-Layer Filter falsely detected checked-in secrets. Their precisions all reached 100%. The high precision of each tool is expected as the Google API key exhibits a strong pattern that can be precisely captured by regular expressions. In addition, both selected tools consider the reduction of false positives to be one of their main objectives. Three-Layer Filter leverages entropy filter, word filter and pattern filter to reduce false positives. LeakScope identifies checked-in secrets via backward slicing starting from the arguments of cloud APIs, ensuring that the extracted strings are checked-in secrets.
However, we made different observations for Twitter client ID/Secret. The 10 Twitter checked-in secrets detected by LeakScope are all considered to be valid Twitter checked-in secrets through manual analysis. In contrast, none of the 383 Twitter checked-in secrets identified by the Three-Layer Filter were considered valid Twitter checked-in secrets. These checked-in secrets are indeed random strings, with some appearing to be tokens for services such as Facebook, CRASHLYTICS, or others. However, none of them are related to Twitter. Such high false positive rates occur because Twitter checked-in secrets are detected with loose regular expressions. This suggests that the effectiveness of the Three-Layer Filter heavily relies on the quality of the regular expressions. For the checked-in secrets that do not exhibit strong patterns, adopting the Three-Layer Filter is likely to result in a large number of false positives.
Exploitable Checked-in Secrets. We acknowledge that our manual analysis may be subject to biases. Our manual analysis process may falsely consider a string to be a valid checked-in secret or vice versa. To address this concern, we validated the Google API keys by sending requests. We implement this validation process to demonstrate the exploitability of the detected checked-in secrets. After validating all the obtained Google API keys by sending a request to the server. 74 Google API keys grant access to the Google Map service. We reported these Google API keys to the developers of the corresponding Android apps, and we received responses from 13 Android apps that acknowledged the issue. This validation signifies that most of the checked-in secrets detected in this study are likely to be exploitable, revealing the severity of the checked-in secrets leakage in Android apps.
Summary: Overall, only 2,142 detected checked-in secrets are certain to be valid after manual validation. Among the tools, LeakScope demonstrated its strength in reducing false positives, as it produced none for both Google API keys and Twitter Client IDs/Secrets. In contrast, while the Three-Layer Filter produced no false positives for Google API keys, it generated a significant number for Twitter Client IDs/Secrets. This indicates that the Three-Layer Filter’s filtering process is less effective when handling loose regular expressions.
4.4. RQ2.3-False Negative
4.4.1. RQ2.3-Methodology
In this research question, we aim to compare the false negatives of each tool. It is impractical to obtain the ground truth of our dataset as it requires manual labelling of over 1 million strings. In addition, all these strings are used in obfuscated code with minimal meaningful contexts. As a result, we cannot calculate the exact recall of each tool. To address this problem, we evaluate the overlap between the detection results of the LeakScope and the Three-Layer Filter. PassFinder was not included as it did not detect any checked-in secrets.
Furthermore, as described in the Section 4.1.3. We could not equalize the detection capability of LeakScope and the Three-Layer Filter. As a result, we focused solely on checked-in secrets categorized as Google API keys for a fair comparison since both of the tools detected a significant number of valid Google API keys. Twitter client ID and secrets are excluded because the Three-Layer Filter failed to detect any valid Twitter client and secrets after manual validation.
4.4.2. RQ2.3-Result
The Venn Diagram in Figure 1 shows the overlap of the detected checked-in secrets between each tool. The Three-Layer Filter identified 374 Google API keys, 11 of which were also detected by LeakScope. Conversely, LeakScope identified 1,816 Google API keys, with 11 overlapping with those found by the Three-Layer Filter. The minimal overlap between the tools suggests that both of them have missed detecting a significant number of checked-in secrets. Even with the absence of the ground truth, this signifies that all the tools have a relatively low recall. However, based on the overlap, we can still conclude that LeakScope missed fewer checked-in secrets compared to others.
We further investigated the checked-in secrets that can only be detected by LeakScope or the Three-Layer Filter. We found that the difference between detection strategies is the main cause of the minimal overlap. For LeakScope, it can extract the strings that are not complete in the Android apps. For example, if a checked-in secret needed to be obtained through several string manipulations such as StringBuilder.append() before being used, the Three-Layer Filter would not be able to detect it as the incomplete string can not match the regular expressions. In contrast, LeakScope is highly dependent on the pre-defined cloud APIs. However, there exist checked-in secrets that are encapsulated in the Android apps but not passed to the cloud APIs. This explains why LeakScope also failed to detect some checked-in secrets. This result suggests that these tools are rather complementary.
Summary: Overall, although the exact number of false negatives could not be determined, the minimal overlap between the tools suggests that both suffer from a significant issue of producing a large number of false negatives.
Figure 1. Venn Diagram for overlap of the number of unique detected Google API keys for LeakScope and Three-Layer Filter.
4.5. A Further Evaluation of PassFinder
As demonstrated in Table 6, PassFinder did not detect any checked-in secrets in this experiment when combining the two models. To thoroughly analyze the performance of PassFinder, we conducted several additional experiments.
4.5.1. Manual Inspection
In our experiment, the password model of PassFinder identified 11,343 strings as potential checked-in secrets. However, the context model concluded that none of these were valid checked-in secrets. To investigate further, we aimed to determine the validity of these strings and explore the reasons behind the context model’s failure to identify any of them. Therefore, we first randomly sampled 372 strings with their contextual surroundings, achieving a 95% confidence level with a 5% margin of error (ran, 2024) from the 11,343 strings identified by the password model. Upon manual investigation, we found that it is similar to the Three-Layer Filter with loose regular expressions; although these 372 strings are randomly generated strings, their context did not provide sufficient information to support that they are valid checked-in secrets. This was primarily due to obfuscation, which significantly differed from the training data used for the context models. The training set consisted of source code extracted from GitHub that contained meaningful, unobfuscated terms. Additionally, the extracted context was presented in Jimple format, an intermediate representation of Java, which further contributed to PassFinder’s ineffectiveness. To evaluate the impact of the Jimple format, we conducted an additional study where the Jimple format was transformed into the Java format. The details of the study are detailed in the following section.
4.5.2. Jimple to Java Transformation
The use of the Jimple format instead of Java may have impacted the performance of the context model. Moreover, since the original sample did not include any valid checked-in secrets, it is challenging to fully determine the effectiveness of the context model. To address this, we created an additional test set containing 664 valid checked-in secrets from Google Cloud services detected by the Three-Layer Filter, along with their contextual surroundings manually converted into Java format. The conversion process was straightforward, as Jimple is an intermediate representation of Java. Listing 4.5.2 shows a Jimple-formatted code snippet from an Android app, closely resembling Java code and providing sufficient information to reconstruct the corresponding Java code. Listing 4.5.2 presents the transformed Java-format snippet from Listing 4.5.2.
{listing}
Extracted Code snippet in Jimple format
{minted}
[…]
4.5.3. Password Model on Labelled dataset
We also conducted an additional experiment to fully validate the performance of the password model as the random sample strings in the previous section did not contain any valid checked-in secrets upon manual verification. Using the same test set from the previous section. This test aimed to evaluate the model’s ability to identify 664 valid checked-in secrets. However, the password model only identified 62 checked-in secrets from the test set, underscoring its restricted capability. The underperformance of the password model may be attributed to the nature of its training set. The machine-generated passwords in the training set were created using a random string generator (Feng et al., 2022), which likely differs from the structure and characteristics of actual checked-in secrets.
5. RQ3: Implications
In RQ3, we discuss the key implications of this research, including the limitations of the existing checked-in secret detection techniques and potential directions for future research in developing a checked-in secret detection tool more suitable for Android apps.
5.1. Limitations of the Evaluated Tools
In this section, we summarize the limitations of the evaluated tools.
5.1.1. Limitation of Three Layer Filter (Intrinsic Value Analysis)
We identify three major limitations of the Three-Layer Filter approach.
First, the detection capability of the Three-Layer Filter is constrained by the manually crafted regular expressions. Regular expression matching is the most important step in the Three-Layer Filter. The number of manually crafted regular expressions determines the detection capability of the Three-Layer Filter. The strings identified by a precise regular expression will likely not require the following filtering process. In our experiment, only one string was excluded from the strings that were obtained through precise regular expressions. However, acquiring precise regular expressions for all categories of checked-in secret can be challenging, and some cloud providers may also generate checked-in secrets that do not exhibit a strong pattern (e.g., Twitter secrets). As a result, it is challenging to expand the detection capability of the Three-Layer Filter.
Second, valid checked-in secrets can be excluded in the arbitrary filtering steps based on entropies, patterns and words. The entropy and pattern filter in our experiment excluded ten strings that are very likely to be a potential checked-in secret through our manual analysis. The threshold of the pattern sequence is set to be a length of four, which is problematic as some checked-in secret could contain such a pattern. For example, the checked-in secret of Firebase Messaging always starts with four repetitive “A”. While the word filter in our experiment did not produce any false negative cases, it’s plausible that checked-in secret could contain English words as one can never make an assumption about the randomly generated string. Furthermore, some checked-in secrets can be stored in a string with a format such as ”key=AIza…”. These checked-in secrets will be excluded as they did not exhibit any regular expression and contain English words such as ”key” and ”token”. As a result, it is hard to define accurate filters to eliminate the noises generated by the regular expressions without excluding false negatives.
Finally, regular expressions cannot identify checked-in secrets that are computed from several strings. After analyzing the checked-in secrets detected exclusively by LeakScope, but which fall within the detection capability of the Three-Layer Filter (same category), we found that the Three-Layer Filter fails to extract checked-in secrets that are obtained through string manipulation methods such as append(). The presence of string manipulation poses a significant limitation for regular expression matching. A critical factor in regular expression matching is the range of the string length. If a checked-in secret is split in the Android apps, it would be impossible for the regular expression to match this string as the length is varied.
5.1.2. Limitation of LeakScope (Static analysis)
We identify four major limitations of LeakScope.
The primary limitation of LeakScope can only detect checked-in secrets from a list of pre-defined cloud APIs. To improve the detection capability of LeakScope, huge manual efforts are needed to expand the list of cloud API signatures. This largely limits the capability of LeakScope.
Second, the unsound static analysis of LeakScope fails to recover some values of the checked-in secrets. LeakScope failed to recover the value of 1,353 identified locations of checked-in secrets. LeakScope is capable of extracting checked-in secrets even if they are stored as environment variables in files such as XML or JSON. However, not all environment variables can be extracted. LeakScope uses APKTool to reverse engineer APK files. Due to the limitations of APKTool, it may sometimes fail to obtain the corresponding files that store environment variables from APK files, leading to the failure of checked-in secret extraction. Moreover, APKTool may also fail to decompile some classes from the APK file, which would affect the performance of LeakScope.
Additionally, since LeakScope uses Soot for backward analysis, it is also subject to the limitations of Soot. Other than the failures caused by the missing files during the reverse engineering process, LeakScope also failed to extract the exact value checked-in secrets that are stored as static variables within a Java class. e.g., public static String Google API key = "Alzaxxxx..". Soot sometimes can only get the name of these variables.
Third, the incorrect Cloud API signature generation and matching can impact the accuracy of detection.. We observed that LeakScope occasionally mismatches the Cloud API in the APK file. Some irrelevant APIs that have the same package structure and similar content would also be considered as one of the pre-defined cloud APIs during the analysis phase of LeakScope.
Finally, there exist checked-in secrets that are not used by cloud APIs. After analyzing the checked-in secrets that were only detected by the Three-Layer Filter, we observed that these checked-in secrets were not used as parameters for any of our selected cloud APIs. Some developers prefer to send requests directly through URLs or use other functions that we are unaware of. As the Venn diagram of our result in Figure 1 indicates, this is not an uncommon situation.
5.1.3. Limitation of PassFinder (Machine-Learning-Based Technique)
When combining the predictions made by both the context model and the password model, PassFinder did not detect any checked-in secrets. To understand the reason behind this, we analyzed the two models separately. We observed that the poor performance of PassFinder can be attributed to (1) the training set and (2) the context modelling under obfuscation and insufficient information from 12 lines of context.
For the password model, the provided training set is a combination of ordinary strings, human-chosen passwords, and machine-generated passwords. These strings may differ from checked-in secrets, although they all appear random. Checked-in secrets of the same cloud service often adhere to specific patterns. As presented in Section 4.5, this difference in characteristics may be the reason why the password model only detected 62 out of 664 checked-in secrets in the training dataset.
For the context model, even though we transformed code snippets from Jimple format to Java format within the training set, the effectiveness of these snippets was diminished by obfuscation. May add some pointers to the previous section where we discussed the further experiments on PassFinder This process replaced variable, method, and class names with random, meaningless letter sequences, reducing their overall significance. Furthermore, we observed that simply 12 lines of code could not always give enough information to detect the presence of checked-in secrets.
5.2. Implications: Future Direction
5.2.1. Complimentary among the tools
As described in Section 4.4, all the tools missed detecting a significant number of checked-in secrets, indicating that the actual recall of each tool is low. This suggests that future research should focus on improving the recall of checked-in secret detection techniques. Given the small overlap between the Three-Layer Filter and LeakScope, it is reasonable to assume that combining these two techniques could enhance the recall of checked-in secret detection.
However, these two techniques share a common limitation: it is difficult to expand their detection capability. The number of regular expressions constrains the Three-Layer Filter’s detection capability, while LeakScope’s detection capability is constrained by the list of pre-defined cloud API signatures. Both regular expressions and cloud APIs are challenging to discover and formalize.
In contrast, machine-learning-based tools, in general, can adapt to a wider range of cases since they do not require any pre-defined patterns. However, our results suggest that simply providing more training sets cannot yield a significant improvement. The effectiveness of the context model is significantly degraded by the presence of obfuscation. Relying on code snippets as input for the model is proven to be ineffective in our evaluation. This finding suggests the need for a new feature extraction approach that is obfuscation-resilient.
5.2.2. String Group Analysis
After analyzing the methods that contain a checked-in secret, we propose to use string groups as features to detect checked-in secrets. A string group refers to all the strings within one single method. In our manual inspection, we observed that string groups that contain checked-in secrets shared some similar patterns. For instance, the names of cloud providers like ”Google” or ”Firebase,” along with terms related to secrets such as ”user,” ”token,” or ”key,” frequently appear in these groups. This suggests that the patterns in string groups could serve as potential indicators of checked-in secrets. Moreover, since string literals are typically not modified by the obfuscation of apps in Google Play Store (Dong et al., 2018).
To verify the feasibility of this approach, we conducted a preliminary study with the hypothesis that string groups containing checked-in secrets exhibit similarities to one another and are distinct from those that do not contain any secrets. To test this hypothesis, we designed a statistical analysis process. First, we collected two collections of string groups: Collection Secret included 602 string groups containing Google API keys and secrets obtained by the Three-Layer Filter. We excluded the string groups that only contain checked-in secret itself. To prevent potential bias in terms of quantity, the second set, Collection NoSecret, included 602 randomly selected string groups that did not contain any checked-in secrets. Methods with only a single string were excluded to ensure meaningful comparison. We then calculated the similarity between string groups using cosine similarity (Han et al., 2012), which is better for this task as it works with non-English text, unlike Levenshtein distance (Navarro, 2001) or Word Movers’ Distance (Saxena, 2019). Cosine similarity measures the cosine of the angle between two character vectors, generating a value between 0 and 1, where 1 indicates identical word vectors and 0 indicates no similarity. The similarity score is calculated using the equation below (Han et al., 2012).
(1) $\text{sim}(x,y)=\frac{x\cdot y}{\|x\|\cdot\|y\|}$
For each string group in Collection Secret, we first transform the string groups into the character vector, in which the number in the vector represents the frequency of a character. Then, we computed the average cosine similarity both within Collection Secret (Comparison-SvS) and between Collection Secret and Collection NoSecret (Comparison-SvN). We need to verify that the two comparisons are statically different to test our hypothesis. We applied both F-test (fte, 2024) and Z-test (zte, 2024). These tests assessed whether the means and variances of the cosine similarity scores in Comparison-SvS and Comparison-SvN were significantly different. The statistical significance was determined through p-values, with smaller p-values indicating a greater likelihood that the differences in means or variances were not due to chance. Additionally, given the various forms that strings can take in Android apps, we modified Collection Secret and Collection NoSecret into three distinct versions: case-sensitive, case-insensitive, and English Word Extraction. The first two variations considered whether string case impacted the similarity calculations, while the English Word Extraction version isolated all English words using a dictionary (dolph, [n. d.]) for further analysis.
The result of the two tests is shown in Table 7. The calculated p values of the two tests for the three versions are all very close to 0 and much smaller than the common significance level alpha (0.01) (alp, 2015), indicating that differences of Comparison S-S and Comparison S-N are statistically significant. We observed that the difference in the case-sensitive comparison is larger than in other variations. This is likely because ordinary strings tend to use more capital letters. Furthermore, the mean similarity score in Comparison S-S is also higher than in Comparison S-N.
Such test results demonstrate the potential to leverage string groups to detect checked-in secrets in obfuscated Android apps. We plan to further explore this direction to propose a new machine-learning based checked-in secret detection technique based on string groups.
Table 7. Result of F-test and Z-test
F-Test(p value) Z-Test(p value) Similarity Score Mean (Comparison s-s: Comparison s-n)
Case-Sensitive $1.76\times 10^{-14}$ 0 0.55 : 0.13
Case-Insensitive $7\times 10^{-13}$ 0 0.65 : 0.54
English Word Extraction $7.6\times 10^{-12}$ 0 0.65 : 0.55
6. Discussion
In this section, we would like to discuss some additional findings and contributions from our research.
Severity of Checked-in Secret Issue in Android apps. Most of the existing studies focused on the checked-in secrets problem on code-sharing platforms, neglecting the realm of Android applications distributed via platforms like the Google Play Store. Consequently, there exists a gap in systematic analyses of checked-in secrets issues within such projects. Our study revealed that excluding the potential false positive, 2142 unique valid checked-in secrets were found among 2,115 Android apps, which accounts for nearly half of our entire dataset. This result underscores the prevalence of the checked-in secret issues in Android apps.
Negligence of Checked-in Secret Issue. We reached out to the developers of Android apps affected by the checked-in secret issue. Some developers acknowledged the problem and committed to addressing it, while others acknowledged the issue but asserted that the checked-in secret was safe to expose. While it is true that some checked-in secrets, such as the Firebase API key, are considered safe to expose in code (though not recommended) (fir, 2024). Or Twitter Client Secret, which would only become harmful if the associated Twitter Client ID is also leaked, these keys can still pose a security threat when other related information is also exposed. For instance, if project IDs and authentication tokens associated with these checked-in secrets are leaked, they can compromise security. In our experiment, we successfully extracted additional information associated with the Firebase API key that could be harmful when combined, demonstrating the potential risk even for seemingly harmless checked-in secrets.
A benchmark dataset for checked-in secrets leakage in Android apps. We have made available the code snippets and the data flow calculated by LeakScope for all the detected checked-in secrets, facilitating future research on this subject. The checked-in secrets and the names of the Android apps are masked due to the sensitivity of the information. However, we are willing to share the complete dataset with researchers and developers upon request.
7. Threats to Validity
First, as described in section 4.3.1 and 4.4.1, the lack of ground truth is a common challenge in checked-in secrets related work. Consequently, a significant threat to our study lies in evaluating the false positives for each tool. Without an available benchmark, it was impossible to determine the exact number of false positives, leaving manual analysis as the only means to verify results. To address this, our manual analysis of detected checked-in secrets would consider not only the format of the strings but also their contextual surroundings. We further validated a part of our checked-in secrets by sending requests to the cloud services, thereby enhancing the rigour of our manual evaluation process. Similarly, we were unable to calculate the recall for each tool. To provide a snapshot of the recall, we analyzed the overlap between the detections of each tool.
Another threat is the detection capability of the selected tools. For instance, the Three-Layer Filter relies on the provided regular expressions as the starting point. Although the authors provided 15 distinct regular expressions, they did not cover all the popular cloud providers. To mitigate this threat, we tried to derive more regular expressions that could find a match in our dataset, and overall, we generated two new regular expressions. In future work, we plan to generate more regular expressions to cover a wider range of cloud providers. As for LeakScope, similar to the Three-Layer Filter, cloud API signatures are the analysis’s starting point. The cloud APIs provided by the authors only cover three cloud providers; we collected 13 more cloud APIs covering five more cloud providers. While our collected cloud APIs cannot cover all the available cloud providers, they did improve the detection capability of LeakScope.
In the end, the training and testing process of PassFinder is also one threat to our work. The training set provided by the authors to train the context model does not contain code snippets in Jimple format, while our provided code snippets are in Jimple format. To mitigate this threat, we transformed the code in Jimple format into JAVA format to truly test the performance of PassFinder.
8. Related work
Checked-in Secret Detection for GitHub. Leakage of checked-in secret issue in the code-sharing platform is a long-studied subject. Besides the tools we selected to run the experiment, there exist other papers that also targeted the checked-in secret issue. Saha et al.  (Saha et al., 2020) introduce a machine learning-based detection tool; they designed 20 features for a string to reduce the false positive, and Lounici et al. (Lounici et al., 2021) examined the performance of various machine learning model when performing checked-in secret detection tasks. There are also some community checked-in secret detection tools. As an example, TruffleHog (tru, 2024), which was originally designed for GitHub, has also been adapted to the Android apps on the Google Play store. However, the Three-Layer Filter, which utilizes the same methodology, outperforms TruffleHog. Some researchers have also proposed tools that can detect human-generated passwords. Fadi et al.  (Sibai et al., 2008) utilized neural networks to predict whether a string is a human-generated password. While Lykousas et al. (Lykousas and Patsakis, 2023) examined the developers’ traits when choosing passwords. Those studies all aimed to develop an effective checked-in secret detection tool for open-source projects but did not examine the feasibility of their approaches on Android apps.
Empircal Study on Checked-in Secrets Issue. There have also been several researches that conducted empirical studies on checked-in secret topics. Basak et al. (Basak et al., 2023a) evaluated the performance of all the available checked-in secret detection tools on a benchmark dataset (Basak et al., 2023b) that contains 818 Github repositories. They also discussed developers’ challenges about checked-in secret by analyzing 779 questions on Stack Exchange (Basak et al., 2023c). Whereas Rahman et al. (Rahman et al., 2022) aimed to investigate the phenomenon of developers disregarding warnings issued by checked-in secret detection tools, often due to the tools’ high false positive rates. Furthermore, Krause et al. (Krause et al., 2022) analyzed the reason why developers would commit checked-in secrets by interviewing other developers. Naiakshina et al.  (Naiakshina et al., 2017) studied how developers deal with checked-in secret storage through an experiment involving 20 developers. However, these empirical analyses have primarily focused on open-source projects, which may not accurately represent the situation in Android apps. We aim to address this limitation in our research.
Security Concerns in Mobile Apps. The popularity of mobile apps has also drawn researchers’ attention. Many studies were conducted to analyze the security issues in mobile apps. PAMDroid (Zhang et al., 2020) was developed to investigate the privacy leakage issue caused by the misconfiguration of analytic service. DroidContext (Wang et al., 2016) targeted the detection of privacy leakage and is capable of differentiating malicious privacy identification and benign privacy disclosure identification. Bunyakiati et al. (Bunyakiati and Sammapun, 2019) examined current secret management and handling in the mobile apps development cycle. Whereas Rahat et al. developed OAUTHLINT (Rahat et al., 2019) to detect the security mistakes of the Android apps that use OAuth API. our work further investigated the security issue caused by checked-in secret.
9. Conclusion
In this paper, we conducted a literature review to categorize the existing checked-in secret detection technique, selected three representative tools and performed an empirical analysis with the selected tools on Android apps after some adaptation. We evaluated the performance of the tools in terms of their effectiveness and efficiency. Overall, 2,142 unique checked-in secrets were detected, affecting 2,115 Android apps. We further performed a systematic comparison among the tools to analyze the limitations of current checked-in secret detection techniques. Based on our findings, we proposed possible research directions for developing a more effective checked-in secret detection tool specifically for Android apps. The checked-in secrets resulted from our experiments form a benchmark for checked-in secret detection in Android apps. The sanitized datasets are available on our project website (dat, [n. d.]) without leaking any sensitive checked-in secrets.